OpenStudy (sammietaygreen):

For which intervals is the function y = x^3+x^2-5x-6 increasing and decreasing?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=3x^2+2x-5\] it is increasing if \[\frac{ dy }{ dx }>0,decreasing~if~\frac{ dy }{ dx }<0\]

OpenStudy (sammietaygreen):

Would it be A??

zepdrix (zepdrix):

Hmm :d

OpenStudy (sammietaygreen):

Am I correct?